It takes Joann 8 hours to do the job alone. Joann and Paulette working together takes 6 hours to do the job. How long will it take Paulette working alone? Give answer in hours and minutes.
Answer by rwm(914)


6/8+6/p=1
It should be obvious that Paulette isn't much help and takes considerably longer to do the job alone since the time was only shortened by 2 hours.
3/4+6/p=1
6/p=1/4
24 hrs=p
check
6/8+6/24==1
3/4+1/4=1
ok
another way to look at
Joan does 1/8 of the job every hour
so in 6 hours she has done 6/8 of the job
which leaves 2/8 for Paulette
[2/8=6/24]
so in 6 hours Paulette has done
6/p=2/8
48=2p
24=p